### v1

@@ Abstract
The current evaluation of Large Language Models (LLMs) predominantly relies on benchmarks focusing on their embedded knowledge by testing through multiple-choice questions (MCQs), a format inherently suited for automated evaluation. Our study extends this evaluation to explore LLMs’ pragmatic competence—a facet previously underexamined before the advent of sophisticated LLMs, specifically in the context of Korean. We employ two distinct evaluation setups: the conventional MCQ format, adapted for automatic evaluation, and Open-Ended Questions (OEQs), assessed by human experts, to examine LLMs’ narrative response capabilities without predefined options. Our findings reveal that GPT-4 excels, scoring 81.11 and 85.69 in the MCQ and OEQ setups, respectively, with HyperCLOVA X, an LLM optimized for Korean, closely following, especially in the OEQ setup, demonstrating a score of 81.56 with a marginal difference of 4.13 points compared to GPT-4. Furthermore, while few-shot learning strategies generally enhance LLM performance, Chain-of-Thought (CoT) prompting introduces a bias toward literal interpretations, hindering accurate pragmatic inference. Considering the growing expectation for LLMs to understand and produce language that aligns with human communicative norms, our findings emphasize the importance for advancing LLMs’ abilities to grasp and convey sophisticated meanings beyond mere literal interpretations.
Index Terms:
Large Language Model (LLM), LLM Evaluation, Pragmatics, Korean.
[…]
=-21pt
I Introduction
@@
Research on LLMs has seen rapid advancements in recent years [1], [2], [3]. Notably, ChatGPT[4], a web-based service built on GPT-3.5[5], was released in November 2022 and exemplifies this remarkable technological advancement. It has demonstrated impressive capabilities in a broad spectrum of Natural Language Processing (NLP) tasks, ranging from traditional ones like sentiment analysis and translation [6], [7] to more demanding areas such as complex problem-solving and creative writing [8], [9]. This versatility of ChatGPT has not only drawn significant attention from NLP researchers but has also captivated the interest of the general public worldwide and numerous tech companies are rigorously entering into the development of their own LLMs [10], [11], [12].
@@
The evaluation of LLMs is as crucial as their development, as it enables the measurement of their performance for the targeted tasks and ensures that these models align with the anticipated standards of capability [13]. Systematic evaluation allows uncovering both the strengths and weaknesses of LLMs, which makes further fine-tuning of the models possible. This cycle of development and evaluation is essential in advancing these models, contributing to the creation of more reliable and suitable LLMs across diverse domains of real-world applications [14], [15].
@@
Benchmarks are serving a significant role in the current evaluation of LLMs [16], [17]. By providing task-specific datasets aligned with corresponding metrics, they establish standardized scenarios in which fine-tuned LLMs can be consistently evaluated. The primary advantage of these benchmarks lies in their ability to automate the evaluation process, allowing a fair comparison of LLMs trained by different developers using varied strategies. However, there are notable limitations in the current benchmarking approach: a predominant focus on aspects like reasoning, computation, and knowledge [18], [19], [20] with an emphasis on literal meaning, rather than implied meanings varying with context, which may not fully capture human-like language understanding. Additionally, many benchmarks rely on MCQs, a format that, while convenient for automated evaluation, does not fully evaluate the generative capacities of LLMs [21]. Furthermore, current benchmarks are showing clear tendency to English-centric evaluation, which results in the under-exploration of LLMs’ multilingual capacities [22], [23].
TABLE I: Variations in Pragmatic Interpretation Based on Context
Statement ”There’s pizza in the fridge.
@@
Pragmatics is a linguistic study dealing with understanding language beyond just the literal meanings of words. It involves interpreting both the explicit (literal) and implicit (nonliteral) aspects of language, heavily influenced by context. Grice, H. P. illuminates this by distinguishing between what is said (literal meaning) and what is implicated (nonliteral meaning), emphasizing the critical role of context in language comprehension [24]. As an illustration, consider Table I, which presents a scenario involving the statement ”There’s pizza in the fridge.” The literal interpretation is straightforward: pizza is inside the fridge. However, the implicated meanings vary with context. In the first scenario, imagine a friend visiting and expressing hunger; the statement might imply permission to eat the pizza. In the second scenario, consider a couple recovering from an argument. Here, the same statement might carry an undertone of reluctance to cook, reflecting the strained atmosphere. These examples illustrate that the pragmatic interpretation of a simple statement can vary significantly, transforming it into complex, context-dependent communication.
@@
While earlier NLP models primarily focused on syntactic and semantic aspects of human language, with little emphasis on pragmatics, current LLMs necessitate a more comprehensive evaluation that extends beyond these traditional aspects [25], [26]. The enhanced performance of these models across diverse NLP tasks marks the significant need for evaluating their contextual language comprehension. Pragmatic understanding is especially crucial for conversational setups where AI assistants are expected to understand and respond in ways that meet human communicative needs [27], [28].
In this paper, we demonstrate a systematic evaluation of LLMs’ pragmatic competence for Korean by analyzing it through four Gricean maxims: quantity, quality, relation, and manner, which are essential for understanding conversational implicature. We parallelize our experiments in two different setups: MCQs to explore LLMs’ pragmatic abilities through automated evaluation and OEQs to assess their narrative generation capacities, thereby revealing biases depending on question types. Furthermore, we investigate the impact of few-shot learning [29] and CoT reasoning [30], in-context learning strategies known to enhance model performance without the need for extensive retraining. By focusing on Korean, we demonstrate LLMs’ capabilities in handling Korean and provide deeper insights into their pragmatic capabilities, including those specifically optimized for Korean. Our work aims to narrow the existing gap between the rapidly evolving capabilities of LLMs and the nuanced, human-level evaluation of language, ultimately contributing to the development of AI systems with a more sophisticated awareness of contextual nuances.
II Related Work
II-A Evaluation Landscape: Methods for Assessing LLMs
The demand for an automated evaluation of language models that combines speed and reliability has become increasingly apparent, aiming to replace the costly and time-consuming process of human assessments. To meet this need, a range of benchmarks have been developed, each designed to evaluate the varied capabilities of different language models with specific focuses [31], [32], [33]. However, as LLMs emerge and evolve, these existing benchmarks have gradually begun to lose their discriminative power, leading to the creation of new benchmarks equipped with more challenging datasets—a task that remains an ongoing process [34], [35], [36]. Today, those benchmarks stand as the primary indicators for claiming the superiority of newly released LLMs.
The qualitative evaluation of generated text is a crucial part of assessing LLMs, as they have evolved as generative language models [37]. Predominantly, current benchmarks are designed to assess the LLMs’ capabilities in language understanding, thus inheriting a notable limitation in a thorough measurement of their generative abilities [18], [19], [20]. In this context, the role of human evaluations becomes indispensable for bridging these existing gaps. By providing a detailed assessment, human evaluation is adept at revealing both the strengths and weaknesses of LLMs at a finer-grained level. Consequently, human evaluation serves to complement aspects that automated evaluations may overlook, and continues to retain a critical importance in the ongoing research of LLMs [38], [39].
Recently, there has been active research into using LLMs as tools for evaluation, a method known as LLM-based evaluation [40], [41], [42]. These evaluation frameworks have shown superior performance compared to other state-of-the-art automated evaluation models, but they still do not reach the level of expert groups and possess limitations such as bias towards long sentences and the text generated by LLMs [42], [43].
II-B Evaluating Pragmatic Competence of LLMs
There have been efforts to assess the pragmatic capabilities of LLMs across varied contexts. Pietro et al. [44] evaluated ChatGPT’s pragmatic ability in Italian using the Test for the Assessment of Pragmatic Abilities and Cognitive Substrates (APACS) [45] as a primary test set, which covers six categories—interview, description, narratives, figurative language 1, 2, and humor. Their findings indicate that, while ChatGPT broadly approximates the pragmatic comprehension capabilities of humans, it has shown a tendency towards over-informativeness. Further challenges are noted in computing text-based inferences, processing physical metaphors, and understanding jokes.
However, the methodology employed raises questions regarding the appropriateness of using human-targeted tests for AI systems. ChatGPT was observed to offer responses considered excessive when addressing questions with inherent human attributes, such as ”Do you live with your family?” or ”What is your house like?”, often by reasserting its AI identity. Viewing these responses as overly informative within this experimental setup overlooks the functional considerations of ChatGPT’s design as an AI assistant, which aims to adhere to AI ethics principles, including self-awareness [46]. Additionally, the lack of transparency concerning the test’s specifics, including the test details, the total number of questions, and the evaluation criteria, undermines the robustness of the experimental evidence.
Bojic et al. [47] evaluated the contextual understanding of implicature in LLMs according to Grice’s Cooperative Principle and its four maxims [24]. The study compared the performance of the selected LLMs—GPT-2, GPT-3, GPT-3.5, GPT-4 and Bard—with that of humans, which was tested in English. They reported that the performance of the LLMs surpassed the human average, with GPT4 showing the highest scores, outperforming all other models and human participants.
Yet, this study is limited in that the human participants were recruited from the engineering faculty of the University of Belgrade in Serbia, specifically those who passed an internally developed English test. This suggests that the evaluation may not reflect the pragmatic competence of native English speakers accurately. Furthermore, the limited number of test items—four per category, culminating in a total of twenty—is deemed insufficient for a representative evaluation of general pragmatic competence, both within each specific category and overall.
II-C Korean-Specific LLMs and its Evaluation
Recently, Korean IT companies have been at the forefront of developing Korean-specific LLMs. In August 2023, NAVER unveiled Clova X, an AI assistant service powered by HyperCLOVA X [48], an LLM specially optimized for Korean. This service is targeting to serve the unique demands of South Korea, aiming to fulfill local service needs. Furthermore, in January 2024, Kakao Brain released Honeybee [49], a Multimodal Large Language Model (MLLM), which integrates language and vision, as open source contributing to further LLM research.
Meanwhile, the development of Korean-specific open-source LLMs is being accelerated and centered around the Open Ko-LLM Leaderboard [50]. This leaderboard challenges LLMs with five benchmarks: four are Korean-adapted versions of existing English benchmarks (ARC [18], HellaSwag [51], MMLU [20], TruthfulQA [52]), and one is the independently developed Ko-CommonGen benchmark [53], which assesses generation capabilities in Korean. As of February 2024, the leaderboard has seen submissions of over thousand LLMs.
Separate from the leaderboard, KMMLU [54] emerges as another critical benchmark for assessing LLMs’ capabilities in Korean. Specifically designed to test performance across 45 diverse categories, it focuses on a broad spectrum of Korean-specific knowledge. This benchmark is anticipated to further facilitate the development of LLMs that accurately embody the region-specific knowledge of Korea.
@@ II-C Korean-Specific LLMs and its Evaluation
While not targeting an LLM, Nam et al. [55] evaluated the pragmatic competence of an AI speaker in Korean, using Gricean maxims in a multi-turn dialogue setup. They found that the maxim of relation was the most frequently violated by the AI speaker. Yet, it is evident that research into the pragmatic abilities of Korean LLMs remains in its initial phases, highlighting both a lack of studies specialized for Korean and the need for more in-depth evaluations of LLMs’ pragmatic understanding.
nanummj
TABLE II: Example of Test Unit on the Maxim of Quality
@@
As Cheol-su was practicing the violin, Yeong-hee, who was listening
next to him, said,
”My dog plays it better.”
MCQ
다음 보기에서 위 발화가 갖는 가장 적절한 의미를 고르세요.
(1) 철수의 바이올린 연주가 형편없다.
(2) 영희가 기르는 강아지는 철수보다 바이올린을 잘 켠다.
(3) 철수의 바이올린 연주는 강아지도 감동시킬 만큼 훌륭하다.
(4) 철수는 고양이를 키우고 있다.
Choose the most appropriate meaning of the statement above from the
options below.
(1) Cheol-su’s violin performance is terrible.
(2) The dog raised by Yeong-hee plays the violin better than Cheol-su.
(3) Cheol-su’s violin performance is so excellent that it can even move
    a dog.
(4) Cheol-su is raising a cat.
OEQ
위 발화가 갖는 가장 적절한 의미를 서술하세요.
Describe the most appropriate meaning of the statement above.
III Methodology
Our study is conducted across three distinct experiments designed to assess the pragmatic competence of LLMs. We begin by evaluating LLMs’ ability to make pragmatic inferences through MCQs, then move to analyze their narrative responses to OEQs to probe their generative capabilities. Lastly, we explore LLMs’ adaptability within in-context learning scenarios, incorporating few-shot learning and CoT prompting.
To further elaborate on our methodology, this section is structured as follows: Section 3.1 introduces Gricean conversational maxims, which form the theoretical foundation for our study. Section 3.2 details the development of a pragmatic test set, crafted specifically to assess the LLMs’ abilities in line with these maxims. Section 3.3 discusses two distinct test setups for MCQs and OEQs, focusing on assessing LLMs’ understanding and response generation. Section 3.4 outlines in-context learning scenarios, detailing our approach to evaluate the effects of few-shot learning and CoT techniques on LLMs’ pragmatic capabilities.
III-A Gricean Conversational Maxims
In evaluating the pragmatic competence of LLMs, we utilize Grice’s theory of conversational implicature as the foundational framework. Implicature refers to the meanings that speakers imply but do not explicitly state, which listeners must deduce from contextual cues. This aspect is essential for effective communication as humans often rely on implied meanings rather than explicit statements in real-world conversations.
Grice delineated the cooperative principle as the foundation for rational conversation. This principle states that participants should make contributions that are appropriate for the current stage of the conversation, guided by its accepted purpose or direction [24]. It is further categorized into four conversational maxims:
  • The Maxim of Quantity: Make your contribution as informative as is required.
  • The Maxim of Quality: Try to make your contribution one that is true.
  • The Maxim of Relation: Ensure that all the information you provide is relevant to the current conversation.
  • The Maxim of Manner: Be perspicuous; Be brief and orderly, and avoid obscurity and ambiguity.
Conversational implicatures are often expressed by intentionally flouting Grice’s maxims, which leads to implicated meanings beyond the literal. A common way to flout the maxim of quantity is through tautology—statements that are redundant or always true, offering no information, such as, “Either the news is true, or it is not true.” Flouting the maxim of quality can involve making obviously false or exaggerated statements, exemplified by, “Sally ran faster than light.” Similarly, flouting the maxim of relation occurs when statements are unrelated to the ongoing conversation, for instance, mentioning the weather during a conversation about exam results—“The weather today is very nice, isn’t it?”—demonstrates this principle. An example of flouting the maxim of manner, which requires clear and orderly expressions, might be a response to a political inquiry at a press conference with, “I would not try to alter your viewpoint about it,” intentionally avoiding a direct answer to maintain ambiguity.
III-B Constructing Pragmatic Test Set
Guided by Grice’s maxims, we have constructed a test set consisting of 120 test units, with 30 units allocated to each maxim. Each test unit is designed to elicit a specific implicated meaning, facilitating an evaluation of LLMs’ capabilities in pragmatic inference. This set was developed collaboratively by four experts holding master’s degrees in linguistics or related fields. Table II showcases an example from our test set, which was originally crafted in Korean and is provided with its English translation. Each test unit consists of a context that sets the scene for the dialogue, a statement made by one of the participants, and a follow-up question that asks the expressed meaning of the statement.

### v2

@@ Abstract
Benchmarks play a significant role in the current evaluation of Large Language Models (LLMs), yet they often overlook the models’ abilities to capture the nuances of human language, primarily focusing on evaluating embedded knowledge and technical skills. To address this gap, our study evaluates how well LLMs understand context-dependent expressions from a pragmatic standpoint, specifically in Korean. We use both Multiple-Choice Questions (MCQs) for automatic evaluation and Open-Ended Questions (OEQs) assessed by human experts. Our results show that GPT-4 leads with scores of 81.11 in MCQs and 85.69 in OEQs, closely followed by HyperCLOVA X. Additionally, while few-shot learning generally improves performance, Chain-of-Thought (CoT) prompting tends to encourage literal interpretations, which may limit effective pragmatic inference. Our findings highlight the need for LLMs to better understand and generate language that reflects human communicative norms. The test set is publicly available on our GitHub repository at https://github.com/DojunPark/pragmatic_eval_korean.
1 Introduction
@@
Research on LLMs has seen rapid advancements in recent years Zhao et al. (2023); Yang et al. (2024). Notably, ChatGPT was released in November 2022 and exemplifies this remarkable technological advancement. It has demonstrated impressive capabilities in a broad spectrum of Natural Language Processing (NLP) tasks, ranging from traditional ones like sentiment analysis and translation Sudirjo et al. (2023); Jiao et al. (2023) to more demanding areas such as complex problem-solving and creative writing Orrù et al. (2023); Shidiq (2023). The versatility of ChatGPT has not only drawn significant attention from NLP researchers but also captivated the general public and tech companies, prompting many to develop their own LLMs Manyika and Hsiao (2023); Touvron et al. (2023).
@@
The evaluation of LLMs is as crucial as their development. It enables the measurement of their performance for the targeted tasks and ensures that these models align with the anticipated standards of capability Chang et al. (2024). Systematic evaluation uncovers both the strengths and weaknesses of LLMs, which makes further fine-tuning of the models possible. This cycle of development and evaluation is essential in advancing these models, contributing to the creation of more reliable and suitable LLMs across diverse domains of real-world applications Lappin (2024); Ge et al. (2024).
@@
Benchmarks are serving a significant role in the current evaluation of LLMs Fourrier et al. (2024). They provide task-specific datasets aligned with metrics, creating standardized scenarios for consistent evaluation. The primary advantage of these benchmarks is automating the evaluation process and enabling fair comparisons of models trained by different developers using varied strategies. However, the current benchmarking approach has notable limitations: a predominant focus on aspects like reasoning, computation, and knowledge (Clark et al., 2018; Hendrycks et al., ; Hendrycks et al., 2020) with an emphasis on literal meaning, rather than implied meanings that vary with context, which are crucial for human-like language understanding. Additionally, many benchmarks rely on MCQs, a format that, while convenient for automated evaluation, does not fully evaluate the generative capacities of LLMs Khatun and Brown (2024). Furthermore, current benchmarks are showing a clear tendency to English-centric evaluation, which results in the under-exploration of LLMs’ multilingual capacities Guo et al. (2023); Bommasani et al. (2023).
Statement "There’s pizza in the fridge."
@@
Table 1: Variations in Pragmatic Interpretation Based on Context
@@
Pragmatics is a linguistic study dealing with understanding language beyond just the literal meanings of words. It involves interpreting both the explicit (literal) and implicit (nonliteral) aspects of language, heavily influenced by context Grice (1975). As an illustration, consider Table 1, which presents a statement with multiple possible interpretations. The literal interpretation is straightforward: pizza is inside the fridge. However, the implicated meanings vary with context. In the first scenario, e.g., imagine a friend visiting and expressing hunger; the statement might imply permission to eat the pizza. In the second scenario, e.g., consider a couple recovering from an argument. Here, the same statement might carry an undertone of reluctance to cook, reflecting the strained atmosphere. These examples illustrate that the pragmatic interpretation of a simple statement can vary significantly, transforming it into complex, context-dependent communication.
@@
While earlier NLP models primarily focused on syntactic and semantic aspects of human language, with little emphasis on pragmatics, current LLMs necessitate a more comprehensive evaluation that extends beyond these traditional aspects Kabbara (2019); Satpute and Agrawal (2022). The enhanced performance of these models across diverse NLP tasks marks the significant need for evaluating their contextual language comprehension. Pragmatic understanding is especially crucial for conversational setups where AI assistants are expected to understand and respond in ways that meet human communicative needs Seals and Shalin (2023a, b).
In this paper, we demonstrate a systematic evaluation of LLMs’ pragmatic competence for the Korean language by analyzing it through four Gricean maxims: quantity, quality, relation, and manner, which are essential for understanding conversational implicature. Through this analysis, we aim to narrow the gap between the rapidly evolving capabilities of LLMs and the nuanced, human-level evaluation of language, ultimately suggesting directions for enhancing AI systems’ awareness of contextual nuances.
Our study provides three main contributions:
  • We introduce the first dedicated resource for pragmatic evaluation of LLMs in Korean.
  • We conduct a comprehensive evaluation of LLMs through MCQ and OEQ setups, assessing both automatic and qualitative dimensions of text generation.
  • We explore the effectiveness of in-context learning strategies, specifically few-shot learning Brown (2020) and CoT reasoning Wei et al. (2022), to demonstrate their potential in enhancing LLM performance.
2 Related Work
Maxim Description
Quantity Make your contribution as informative as is required.
Quality Try to make your contribution one that is true.
Relation Ensure that all the information you provide is relevant to the current conversation.
Manner Be perspicuous; Be brief and orderly, and avoid obscurity and ambiguity.
Table 2: Gricean Maxims of Conversational Implicature
2.1 Gricean Conversational Maxims
In pragmatics, implicature refers to the meanings that speakers imply but do not explicitly state, which listeners must deduce from contextual cues. This aspect is essential for effective communication as humans often rely on implied meanings rather than explicit statements in real-world conversations.
Grice outlined the cooperative principle as the foundation for rational conversation. This principle states that participants should make contributions that are appropriate for the current stage of the conversation, guided by its accepted purpose or direction Grice (1975). This principle is further categorized into four conversational maxims, as demonstrated in Table 2, which are crucial for understanding implicated meanings in communication. Conversational implicatures are often expressed by intentionally flouting these maxims, which leads to implicated meanings beyond the literal.
2.2 Evaluating Pragmatic Competence of LLMs
There have been efforts to assess the pragmatic capabilities of LLMs. di San Pietro et al. (2023) assessed ChatGPT’s pragmatic skills in Italian with the APACS Test Arcara and Bambini (2016), focusing on categories such as figurative language, humor, and interviews. Their results show that although ChatGPT closely mirrors human pragmatic understanding, it tends to be overly informative and struggles with text-based inferences, physical metaphors, and humor comprehension. However, The lack of transparency regarding the full test set limits how their findings align with further research.
Bojic et al. (2023) evaluated LLMs against Grice’s Cooperative Principle and its four maxims, reporting that the LLMs’ performance exceeded the human average, with GPT-4 scoring the highest. This study, however, was limited by a participant pool of non-native English speakers and a small number of test items (twenty total). These factors may not accurately reflect native English speakers’ pragmatic competence, suggesting a need for a larger, publicly available test set for more reliable evaluations of LLMs.
2.3 Korean-Specific LLMs and its Evaluation
The development of Korean-specific open-source LLMs has been accelerated by the introduction of the Open Ko-LLM Leaderboard Park et al. (2024), which features five benchmarks. KMMLU Son et al. (2024) also emerges as an important benchmark, specifically designed to evaluate LLMs’ capabilities in Korean across 45 diverse categories.
@@ 2.3 Korean-Specific LLMs and its Evaluation
While not targeting an LLM, Nam et al. Nam et al. (2023) evaluated the pragmatic competence of an AI speaker in Korean, using Gricean maxims in a multi-turn dialogue setup. They found that the maxim of relation was the most frequently violated by the AI speaker. Despite these efforts, research into the pragmatic abilities of LLMs for Korean is still in its early stages, underscoring the need for more specialized studies of these LLMs’ pragmatic understanding.
{CJK}
[…]
@@
Context & Statement 철수가 바이올린을 연습하자 옆에서 듣던 영희가 말 했다. "우리집 강아지가 더 잘 한다." As Cheol-su was practicing the violin, Yeong-hee, who was listening next to him, said, "My dog plays it better."
MCQ 다음 보기에서 위 발화가 갖는 가장 적절한 의미를 고르세요. (1) 철수의 바이올린 연주가 형편없다. (2) 영희가 기르는 강아지는 철수보다 바이올린을 잘        켠다. (3) 철수의 바이올린 연주는 강아지도 감동시킬 만큼        훌륭하다. (4) 철수는 고양이를 키우고 있다. Choose the most appropriate meaning of the statement above from the options below. (1) Cheol-su’s violin performance is terrible. (2) The dog raised by Yeong-hee plays the violin better        than Cheol-su. (3) Cheol-su’s violin performance is so excellent that        it can even move a dog. (4) Cheol-su is raising a cat.
OEQ 위 발화가 갖는 가장 적절한 의미를 서술하세요. Describe the most appropriate meaning of the statement above.
Table 3: Example of a Test Unit on the Maxim of Quality. The answer (1) in bold is the correct answer, as it accurately conveys the implicated meaning of the statement within the provided context.
3 Methodology
3.1 Constructing Pragmatic Test Set
The development of the pragmatic test set was planned to thoroughly assess the nuanced understanding of conversational implicatures by LLMs. Below are the detailed considerations involved in the test set construction:
  • Selection of Maxims: We chose Grice’s maxims as the foundational framework due to their comprehensive coverage of conversational implicatures. These maxims are essential for assessing a model’s ability to interpret meanings beyond literal words.
  • Test Set Size and Distribution: The test set comprises 120 units, with 30 units allocated to each of the four maxims. This distribution ensures a balanced assessment across different aspects of pragmatic competence while allowing for statistically significant results.
  • Contextual Design: Each test unit consists of a context that sets the scene for the dialogue, a statement made by one of the dialogue participants, and a follow-up question that asks the expressed meaning of the statement.
  • Expert Collaboration: The test units were crafted by four experts holding master’s degrees in linguistics or related fields, ensuring high-quality and contextually rich scenarios.
Table 3 presents an example from our test set, demonstrating the case of the maxim of quality. The example is shown in both Korean and its English translation. In this instance, the statement by Yeong-hee, “My dog plays it better,” may seem to simply praise the dog’s abilities if taken at face value. However, within the provided context of Cheol-su practicing the violin, it implies a critical judgment, suggesting that Cheol-su’s violin playing is exceptionally poor.